Computech Corporation is expanding rapidly and currently needs to retain all of its earnings; hence, it does not pay dividends. However, investors expect Computech to begin paying dividends, beginning with a dividend of $2.00 coming 3 years from today. The dividend should grow rapidly - at a rate of 17% per year - during Years 4 and 5, but after Year 5, growth should be a constant 10% per year. If the required return on Computech is 13%, what is the value of the stock today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

D3=2

D4=(2*1.17)=2.34

D5=(2.34*1.17)=2.7378

Value after year 5=(D5*Growth rate)/(Required return-Growth rate)

=(2.7378*1.1)/(0.13-0.1)

=100.386

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2/1.13^3+2.34/1.13^4+2.7378/1.13^5+100.386/1.13^5

=$58.79(Approx)